OpenStudy (anonymous):

At the end of a 5 year term what is the balance of an 8000 investment at a rate of 1.8 percent compound quarterly? round to nearest dollar 8752 8830 8924 8851 everytime i try to do this i dont even get one of the answers help me please.. compound quarterly means 4 times a year

jimthompson5910 (jim_thompson5910):

Use the formula A = P(1+r/n)^(n*t)

jimthompson5910 (jim_thompson5910):

In this case P = 8000 r = 0.018 n = 4 t = 5

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

i tried thats the formula i have been using i keep getting outrageous numbers

jimthompson5910 (jim_thompson5910):

sry the site crashed on me

jimthompson5910 (jim_thompson5910):

this is what you should get A = P(1+r/n)^(n*t) A = 8000(1+0.018/4)^(4*5) A = 8000(1+0.0045)^(20) A = 8000(1.0045)^(20) A = 8000(1.09395339818682) A = 8751.62718549459 A = 8752

OpenStudy (anonymous):

yea it just let me back on. well i kept getting like really long numbers.. i wonder what i did wrong. thank you

jimthompson5910 (jim_thompson5910):

hmm look through my steps and try to replicate them where you differ, tell me and we'll try to see what goes haywire